# Corresponding author(s) Equal contribution

### v1

@@ Abstract
DeepSeek-R1, known for its low training cost and exceptional reasoning capabilities, has achieved state-of-the-art performance on various benchmarks. However, detailed evaluations from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs. To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, and DeepSeek-R1-Distill-Llama series on A-Eval, an application-driven benchmark. By comparing original instruction-tuned models with their distilled counterparts, we analyze how reasoning enhancements impact performance across diverse practical tasks. Our results show that reasoning-enhanced models, while generally powerful, do not universally outperform across all tasks, with performance gains varying significantly across tasks and models. To further assist users in model selection, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models, ensuring optimal performance and resource efficiency in real-world applications.
@@ 1 Introduction
To address the gap, we evaluate DeepSeek models on A-Eval [9], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories. Through this evaluation, we extend existing evaluations and provide actionable insights into how reasoning enhancements influence model performance across various practical tasks. These results and findings help guide users in selecting the most suitable DeepSeek models for their specific needs, enabling cost-effective deployment in real-world applications.
@@ 2.1 Evaluated Models
Llama-3.3-70B-Instruct Dense model SFT from Llama-3.3-7B-Base
@@ 2.2 Dataset
The A-Eval benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across five major categories and 27 subcategories. We use A-Eval’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications. This evaluation helps assess how DeepSeek-R1’s reasoning enhancements improve model capabilities in different practical scenarios.
@@ 2.3 Evaluation Process
We follow A-Eval’s zero-shot automatic evaluation process, which includes:
@@ 3.1 Overall Performance
(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 178.74% and 54.36%, respectively. This is because they are math-focused and perform poorly on general tasks, but the distilled reasoning data from DeepSeek-R1 significantly enhances their general capabilities.
Figure 1: Average score of DeepSeek series models on A-Eval. (a) The overall average score of models across all data. (b) to (f). The average scores of models on each task. The "Instruct Model" refers to the original models without reasoning enhancement, and the "Reasoning Enhancement Model" refers to models that have been enhanced using DeepSeek-R1’s reasoning data.
@@
(5) Before and after distillation, Qwen2.5-32B outperforms the larger model, Llama-3.3-70B.
@@ 3.2 Performance Comparison By Task
(7) The two mathematical models achieve comparable performance in Logical Reasoning but under-perform in other tasks.
@@
(9) Performance degradation occurs after distillation in: Text Understanding (Qwen2.5-14B, Qwen2.5-32B, and DeepSeek-V3), Information Extraction (Llama-3.1-8B, Qwen2.5-14B), and Text Generation (Qwen2.5-14B, DeepSeek-V3). For Task Planning, apart from the two mathematical models, other models either remain unchanged or see a decline in performance after distillation.
@@ 3.3 Performance Comparison By Subtask
(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-11.3%), Part-of-Speech Tagging (-13.4%), and Open-source Question Answering (-17.05%).
@@
(15) For arithmetic operations tasks, distillation provides benefits to all models except for Qwen2.5-Math-7B (85.26 to 84).
@@ 3.5 Model Selection Guidance for Users
Table 2: Performance tier classification of DeepSeek models across major task categories. Models are categorized into tiers (A+: >85, A: 80-85, B: 70-80, C: 60-70, D: <60) based on their scores in Text Understanding (TU), Information Extraction (IE), Text Generation (TG), Logical Reasoning (LR), and Task Planning (TP).
@@
DeepSeek-R1-Distill-Qwen-1.5B 1.5B D D D C D
Qwen2.5-Math-7B-Instruct 7B D D D C D
DeepSeek-R1-Distill-Qwen-7B 7B C D C B B
Llama-3.1-8B-Instruct 8B C C C D B
DeepSeek-R1-Distill-Llama-8B 8B C C B C B
Qwen2.5-14B-Instruct 14B B B A B A+
DeepSeek-R1-Distill-Qwen-14B 14B B B B A A
Qwen2.5-32B-Instruct 32B B B A B A
DeepSeek-R1-Distill-Qwen-32B 32B B B A B A+
Llama-3.3-70B-Instruct 70B B B B B A
DeepSeek-R1-Distill-Llama-70B 70B B B B A A
@@
DeepSeek-R1 671B|37B A A A A+ A
To further simplify the model selection process for users without extensive knowledge of LLMs, we classify the performance of DeepSeek models into five tiers (A+, A, B, C, D) based on their scores in the five major task categories. As shown in Tab. 2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 60). This tiered classification provides an intuitive and accessible way for users to quickly identify the most suitable models for their specific application needs, balancing performance and cost-effectiveness.
@@
Task Planning: Qwen2.5-14B-Instruct (A+)
@@
Text Generation: DeepSeek-R1-Distill-Llama-8B (B)
@@
Task Planning: DeepSeek-R1-Distill-Qwen-7B (B)
@@ 4 Conclusion
To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models and their distilled variants on the A-Eval benchmark. Our analysis reveals that reasoning-enhanced models, while generally powerful, are not universally superior across all tasks. Finally, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models based on their specific application requirements. In future work, we plan to expand the tasks and datasets of A-Eval to offer even more comprehensive insights for model selection and application.

### v2

@@ Abstract
DeepSeek-R1, known for its low training cost and exceptional reasoning capabilities, has achieved state-of-the-art performance on various benchmarks. However, detailed evaluations from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs. To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, and DeepSeek-R1-Distill-Llama series on the improved version A-Eval (A-Eval-2.0), an application-driven benchmark. By comparing original instruction-tuned models with their distilled counterparts, we analyze how reasoning enhancements impact performance across diverse practical tasks. Our results show that reasoning-enhanced models, while generally powerful, do not universally outperform across all tasks, with performance gains varying significantly across tasks and models. To further assist users in model selection, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models, ensuring optimal performance and resource efficiency in real-world applications. It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results. We will continuously optimize the evaluation benchmarks and periodically update this paper to provide more comprehensive and accurate evaluation results. Please refer to the latest version of the paper for the most recent results and conclusions.
@@ 1 Introduction
To address the gap, we first create A-Eval-2.0 by modifying some QAs in A-Eval [9], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories. Then, we evaluate DeepSeek models on A-Eval-2.0. Through this evaluation, we extend existing evaluations and provide actionable insights into how reasoning enhancements influence model performance across various practical tasks. These results and findings help guide users in selecting the most suitable DeepSeek models for their specific needs, enabling cost-effective deployment in real-world applications.
It is important to emphasize that the selection of test samples and the design of evaluation criteria will inevitably introduce certain biases into the evaluation results. To address this, we will continue to optimize this evaluation work to enhance its comprehensiveness and reliability as much as possible.
@@ 2.1 Evaluated Models
Llama-3.3-70B-Instruct Dense model SFT from Llama-3.3-70B-Base
@@ 2.2 Dataset
The A-Eval benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across five major categories and 27 subcategories. To accommodate the enhanced reasoning capabilities of the DeepSeek series models, we refine the QAs of Task Planning category in A-Eval, thereby constructing the evaluation dataset A-Eval-2.0 utilized in this evaluation. We use A-Eval-2.0’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications. This evaluation helps assess how DeepSeek-R1’s reasoning enhancements improve model capabilities in different practical scenarios.
@@ 2.3 Evaluation Process
We follow the automatic and manual scoring methods of A-Eval, and the automatic scoring process includes:
@@ 3.1 Overall Performance
(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 169% and 53.6%, respectively. This is because they are math-focused and perform poorly on general tasks, but the distilled reasoning data from DeepSeek-R1 significantly enhances their general capabilities.
Figure 1: Average score of DeepSeek series models on A-Eval-2.0. (a) The overall average score of models across all data. (b) to (f). The average scores of models on each task. The "Instruct Model" refers to the original models without reasoning enhancement, and the "Reasoning Enhancement Model" refers to models that have been enhanced using DeepSeek-R1’s reasoning data.
@@
(5) Before and after distillation, Qwen2.5-32B outperforms the larger model, Llama-3.3-70B. This may be attributed to the fact that our evaluation dataset includes many Chinese data.
@@ 3.2 Performance Comparison By Task
(7) The two mathematical models achieve comparable performance in Logical Reasoning and Task Planning but under-perform in other tasks.
@@
(9) Performance degradation occurs after distillation in: Text Understanding (Qwen2.5-14B, Qwen2.5-32B, and DeepSeek-V3), Information Extraction (Llama-3.1-8B, Qwen2.5-14B), and Text Generation (Qwen2.5-14B, DeepSeek-V3), and Task Planning (Llama-3.1-8B, Llama-3.3-70B).
@@ 3.3 Performance Comparison By Subtask
(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-9.37), Part-of-Speech Tagging (-10.62), and Open-source Question Answering (-12.38%).
@@ 3.5 Model Selection Guidance for Users
Table 2: Performance tier classification of DeepSeek models across major task categories. Models are categorized into tiers (A+: >85, A: 80-85, B: 75-80, C: 65-75, D: <65) based on their scores in Text Understanding (TU), Information Extraction (IE), Text Generation (TG), Logical Reasoning (LR), and Task Planning (TP).
@@
DeepSeek-R1-Distill-Qwen-1.5B 1.5B D D D D C
Qwen2.5-Math-7B-Instruct 7B D D D D D
DeepSeek-R1-Distill-Qwen-7B 7B C D D B C
Llama-3.1-8B-Instruct 8B C D C D B
DeepSeek-R1-Distill-Llama-8B 8B C D C C C
Qwen2.5-14B-Instruct 14B B B A B C
DeepSeek-R1-Distill-Qwen-14B 14B B C B A B
Qwen2.5-32B-Instruct 32B B B A B C
DeepSeek-R1-Distill-Qwen-32B 32B B B A B A
Llama-3.3-70B-Instruct 70B C C C C B
DeepSeek-R1-Distill-Llama-70B 70B B C B A C
@@
DeepSeek-R1 671B|37B A A A A+ A+
To further simplify the model selection process for users without extensive knowledge of LLMs, we classify the performance of DeepSeek models into five tiers (A+, A, B, C, D) based on their scores in the five major task categories. As shown in Tab. 2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 65). This tiered classification provides an intuitive and accessible way for users to quickly identify the most suitable models for their specific application needs, balancing performance and cost-effectiveness.
@@
If users require A+ level performance, the following models are recommended for each task category:
Text Understanding: DeepSeek-V3 (A) or DeepSeek-R1 (A)
Information Extraction: DeepSeek-R1 (A)
Text Generation: DeepSeek-V3 (A+)
Logical Reasoning: DeepSeek-R1 (A+)
Task Planning: DeepSeek-V3 (A+) or DeepSeek-R1 (A+)
@@
Task Planning: DeepSeek-R1-Distill-Qwen-32B (A)
@@
Text Generation: Qwen2.5-14B-Instruct (A)
@@
Task Planning: DeepSeek-R1-Distill-Llama-8B (B)
@@
It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria. Therefore, our evaluation results may not be the most objective, comprehensive, or accurate. As we will continue to improve this evaluation work, we recommend referring to the latest version of the paper for the most recent experimental results and conclusions.
@@ 4 Conclusion
To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models and their distilled variants on the modified A-Eval-2.0 benchmark. Our analysis reveals that reasoning-enhanced models, while generally powerful, are not universally superior across all tasks. Finally, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models based on their specific application requirements. In the future, we will continue to advance this work by optimizing evaluation benchmarks and promptly updating evaluation results to the community.